OpenStudy (superhelp101):

Solve for x. 7^(x-2)=12^(x)

OpenStudy (juanpablojr):

7^(x-2)=12^(x) ln(7^(x-2))=ln(12^(x)) xln(7)-2ln(7)=ln(12^(x)) xln(7)-2ln(7)-xln(12)=0 x(ln(7)-ln(12))=2ln(7) x=(2ln(7))/(ln(7)-ln(12)) x=3.8918

OpenStudy (juanpablojr):

i double checked.

OpenStudy (superhelp101):

I lol I saw that on another website, but I don't think it's right :P

OpenStudy (superhelp101):

I checked it doesn't work :/ @juanpabloJR

OpenStudy (juanpablojr):

hold on